OpenStudy (anonymous):

If R1 = {(a,b)|a=b(mod3)} and R2 = {(a,b)|a=b(mod4)} What would be the union of these two?

OpenStudy (anonymous):

nvm I think I got it: {(a,b) | a-b=4,8,3,6, or 9 (mod 12)} I think. Not entirely sure

OpenStudy (anonymous):

for R1: a-b=3,6,9,... for R2: a-b=4,8,12,... the union should be a-b=3,4,6,8,9,12,... i don't have any idea how to represent these numbers in modulo :(

OpenStudy (anonymous):

i don't think mod 12 represents the union of R1 and R2. it represents the intersection of R1 and R2.